# Palm distributions of superposed point processes for statistical inference

### v1

@@ Acknowledgments
The third lemma describes the $k$-th order factorial moment measure of a sncp. To state this result in a rigorous way, let us introduce the space $\mathcal{T}_{k}$ as follows. Let $\bm{t}=(t_{1},\ldots,t_{k})$ be a vector of natural numbers such that $\max_{j}t_{j}=|\bm{t}|$, where $|\bm{t}|$ denotes the number of distinct values in $\bm{t}$. Consider the partition $\mathcal{P}_{k}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{|\bm{t}|}\}$ of $\{1,\ldots,k\}$ induced by the ties in $\bm{t}$, such that $j\in\mathcal{C}_{h}$ iff $t_{j}=h$. Then, the space $\mathcal{T}_{k}$ contains the equivalence classes of $\bm{t}$ inducing the same partition $\mathcal{P}_{k}$. This is the correct space where the vector of latent indicators $\bm{T}$ takes value in Theorem 2 and Theorem 3.
@@
$M^{(k)}_{\Phi}(\mathrm{d}\bm{x})=\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\eta(\bm{x}_{h})\mathrm{d}\bm{x},$
where $\bm{x}_{h}=(x_{j}:t_{j}=h)$, for $h=1,\ldots,|\bm{t}|$.
@@
The last theorem characterizes the joint distribution of the number of points $N$ of $\Phi$, when $\Phi$ is a sncp, together with the latent partition among its points, determined by the $\Phi_{i}$’s (see Section 4.1). This result turns out to be useful for proving Theorem 2.
@@
$\mathsf{P}(N,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})=\frac{1}{N!}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\nu(\mathrm{d}\theta\,\mathrm{d}\gamma),$
@@
S4.2 Proof of Theorem 2: the Palm distributions of shot noise Cox processes
To prove Theorem 2 we proceed by induction. For $k=1$ the result is given in Mo03Cox and for $k=2$ the result can be shown by direct calculation thanks to the Palm algebra. Suppose now the statements hold for a general $k$-tuple $\bm{x}_{k}=(x_{1},\ldots,x_{k})$ of distinct points, and let $x_{k+1}\neq x_{j}$, $j=1,\ldots,k$. By the Palm algebra
$\Phi^{!}_{\bm{x}_{k+1}}=(\Phi^{!}_{\bm{x}_{k}})_{x_{k+1}}.$
Observe that $\Phi^{!}_{\bm{x}_{k}}$ has a mixture representation (by the induction hypothesis). Let $\bm{t}_{k}=(t_{1},\ldots,t_{k})\in\mathcal{T}_{k}$ be indicators describing a partition of $\bm{x}_{k}$ into $|\bm{t}|$ clusters. For ease of notation, define $\Psi_{\bm{t}_{k}}=\Phi+\sum_{h=1}^{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}$ where the $\Phi_{\zeta_{\bm{x}_{h}}}$’s are as in the statement of the theorem. Then
$\mathbf{P}_{\Phi^{!}_{\bm{x}_{k}}}=\sum_{\bm{t}_{k}\in\mathcal{T}_{k}}\mathsf{P}(\bm{T}_{k}=\bm{t}_{k})\mathbf{P}_{\Psi_{\bm{t}_{k}}}.$
An application of Proposition S1 yields
$\mathbf{P}_{\Phi^{!}_{\bm{x}_{k+1}}}=\sum_{\bm{t}_{k}\in\mathcal{T}_{k}}w_{\bm{t}_{k}}(x_{k+1})\mathbf{P}^{!x_{k+1}}_{\Psi_{\bm{t}_{k}}},$
where $\mathbf{P}^{!x_{k+1}}_{\Psi_{\bm{t}_{k}}}$ is the law of the process $\left(\Phi+\sum_{h=1}^{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}\right)^{!}_{x_{k+1}}$ and, by Lemma S2,
$w_{\bm{t}_{k}}(x_{k+1})\propto\mathsf{P}(\bm{T}_{k}=\bm{t}_{k})\left(\eta(x_{k+1})+\sum_{h=1}^{\bm{t}_{k}}\frac{\eta(\bm{x}_{h},x_{k+1})}{\eta(\bm{x}_{k})}\right)=:\mathsf{P}(\bm{T}_{k}=\bm{t}_{k})\lambda_{\bm{t}}(x_{k+1}).$
@@
$\left(\Phi+\sum_{h=1}^{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}\right)^{!}_{x_{k+1}}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\begin{cases}\Phi+\Phi_{\zeta_{x_{k+1}}}+\sum_{h=1}^{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}&\text{with prob. }\frac{\eta(x_{k+1})}{\lambda_{\bm{t}}(x_{k+1})}\\ \Phi+\Phi_{\zeta_{(\bm{x}_{j},x_{k+1})}}+\sum_{h\neq j}{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}&\text{with prob. }\frac{\eta(\bm{x}_{h},x_{k+1})/\eta(\bm{x}_{h})}{\lambda_{\bm{t}}(x_{k+1})}\end{cases}$
@@
$\Phi^{!}_{\bm{x}_{k+1}}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\begin{cases}\Phi+\Phi_{\zeta_{x_{k+1}}}+\sum_{h=1}^{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}&\text{with prob. }w_{\bm{t}_{k}}(x_{k+1})\frac{\eta(x_{k+1})}{\lambda_{\bm{t}}(x_{k+1})}\\ \Phi+\Phi_{\zeta_{(\bm{x}_{j},x_{k+1})}}+\sum_{h\neq j}{|\bm{t}_{k}|}\Phi_{\zeta_{\bm{x}_{h}}}&\text{with prob. }w_{\bm{t}_{k}}(x_{k+1})\frac{\eta(\bm{x}_{h},x_{k+1})/\eta(\bm{x}_{h})}{\lambda_{\bm{t}}(x_{k+1})}.\end{cases}$
The proof follows by observing that $w_{\bm{t}_{k}}(x_{k+1})\frac{\eta(x_{k+1})}{\lambda_{\bm{t}}(x_{k+1})}\propto\prod_{h=1}^{|\bm{t}_{k}|}\eta(\bm{x}_{h})\eta(x_{k+1})$ and $w_{\bm{t}_{k}}(x_{k+1})\allowbreak\frac{\eta(\bm{x}_{h},x_{k+1})/\eta(\bm{x}_{h})}{\lambda_{\bm{t}}(x_{k+1})}\propto\prod_{j\neq h}^{|\bm{t}|}\eta(\bm{x}_{j})\eta(\bm{x}_{h},x_{k+1})$.
S4.3 General statement and proof of Theorem 3: the Janossy measures of shot noise Cox processes
We provide the Janossy measures for finite sncps, for any arbitrary intensity measure $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)$. Theorem 3 corresponds to the special case with the additional assumption $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)=\rho(\mathrm{d}\gamma)G_{0}(\mathrm{d}\theta)$. This special case is discussed in the following statement as well.
@@
$j_{k}(x_{1},\ldots,x_{k})=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma).$
@@
$j_{k}(x_{1},\ldots,x_{k})=k!\,\mathsf{P}(\Phi(\mathbb{X})=k)\,\mathsf{E}\left[\prod_{h=1}^{|\bm{T}|}\int_{\mathbb{X}}\prod_{j:T_{j}=h}\kappa(x_{j};\theta)G_{0}(\mathrm{d}\theta)\right],$
where the expectation is taken with respect to the indicators $\bm{T}=(T_{1},\ldots,T_{k})\in\mathcal{T}_{k}$ with distribution
$\mathsf{P}(\bm{T}=\bm{t})\propto\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma).$
@@
The proof follows from specializing Theorem S1 for a sncp. In particular, Theorem S1 states that
$J_{k}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k})=\mathsf{P}(\Phi^{!}_{(x_{1},\ldots,x_{k})}(\mathbb{X})=0)M^{(k)}_{\Phi}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}).$ (S9)
Focusing on the first term of (S9), and denoting with $\bm{x}=(x_{1},\ldots,x_{k})$,
$\mathsf{P}\left(\Phi^{!}_{\bm{x}}(\mathbb{X})=0\right)=\mathsf{E}\left[\mathsf{P}\left(\Phi^{!}_{\bm{x}}(\mathbb{X})=0\,|\,\bm{T}\right)\right]=\mathsf{E}\left[\mathsf{P}\left(\Phi(\mathbb{X})=0\right)\prod_{h=1}^{|\bm{T}|}\mathsf{P}\left(\Phi_{\zeta_{\bm{x}_{h}}}(\mathbb{X})=0\right)\right],$ (S10)
where $\bm{T}$ and the $\Phi_{\zeta_{\bm{x}_{h}}}$’s are introduced in Theorem 2. It follows that (S10) writes as
$\mathsf{P}\left(\Phi^{!}_{\bm{x}}(\mathbb{X})=0\right)$ $=\mathsf{E}\left[\mathsf{P}\left(\Phi(\mathbb{X})=0\,|\,\Lambda\right)\right]\mathsf{E}\left[\prod_{h=1}^{|\bm{T}|}\mathsf{E}\left[\mathsf{P}\left(\Phi_{\zeta_{\bm{x}_{h}}}(\mathbb{X})=0\,|\,\zeta_{\bm{x}_{h}}\right)\right]\right]$
$=\mathsf{E}\left[e^{-\int_{\mathbb{X}}\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(x;\theta)\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}x}\right]\mathsf{E}\left[\prod_{h=1}^{|\bm{T}|}\mathsf{E}\left[e^{-\int_{\mathbb{X}}\gamma_{\bm{x}_{h}}\kappa(x;\theta_{\bm{x}_{h}})\mathrm{d}x}\right]\right]$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\mathsf{E}\left[\prod_{h=1}^{|\bm{T}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\bm{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\right]$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\bm{t}\in\mathcal{T}_{k}}\left[\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\bm{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\right]\mathsf{P}(\bm{T}=\bm{t}),$
where $f_{\bm{x}_{h}}$ is defined in Theorem 2. Finally, plugging this last expression in (S9) and using Lemma S3 for $M^{(k)}_{\Phi}$, we obtain
$J_{k}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k})$ $=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\eta(\bm{x}_{h})\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\bm{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\;$
$\hskip 256.0748pt\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k},$
@@
$\begin{split}&j_{k}(x_{1},\ldots,x_{k})\\ \quad&=\exp\left\{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)\right\}\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{X}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)G_{0}(\mathrm{d}\theta)\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma).\end{split}$ (S11)
Specializing the result in Theorem S2 for $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)=\rho(\mathrm{d}\gamma)G_{0}(\mathrm{d}\theta)$, we obtain
$\mathsf{P}(N,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})=\frac{1}{N!}e^{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma),$
@@
$\mathsf{P}(N=k)$ $=\sum_{\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\}}\mathsf{P}(N=k,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})$
@@
$=\frac{1}{k!}e^{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)}\sum_{\bm{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\bm{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma),$

### v2

@@ Acknowledgments
The third lemma describes the $k$-th order factorial moment measure of a sncp. To state this result in a rigorous way, let us introduce the space $\mathcal{T}_{k}$ as follows. Let $\underline{t}=(t_{1},\ldots,t_{k})$ be a vector of natural numbers such that $\max_{j}t_{j}=|\underline{t}|$, where $|\underline{t}|$ denotes the number of distinct values in $\underline{t}$. Consider the partition $\mathcal{P}_{k}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{|\underline{t}|}\}$ of $\{1,\ldots,k\}$ induced by the ties in $\underline{t}$, such that $j\in\mathcal{C}_{h}$ iff $t_{j}=h$. Then, the space $\mathcal{T}_{k}$ contains the equivalence classes of $\underline{t}$ inducing the same partition $\mathcal{P}_{k}$. This is the correct space where the vector of latent indicators $\underline{T}$ takes value in Theorem˜3 and Theorem˜4.
@@
$M^{(k)}_{\Phi}(\mathrm{d}\underline{x})=\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\eta(\underline{x}_{h})\mathrm{d}\underline{x},$
where $\underline{x}_{h}=(x_{j}:t_{j}=h)$, for $h=1,\ldots,|\underline{t}|$.
@@
The last theorem characterizes the joint distribution of the number of points $N$ of $\Phi$, when $\Phi$ is a sncp, together with the latent partition among its points, determined by the $\Phi_{i}$’s (see Section 4). This result turns out to be useful for proving Theorem˜3.
@@
${\rm pr}(N,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})=\frac{1}{N!}e^{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\nu(\mathrm{d}\theta\,\mathrm{d}\gamma),$
@@
S6.2 Proof of Theorem˜3: the Palm distributions of shot noise Cox processes
To prove Theorem˜3 we proceed by induction. For $k=1$ the result is given in Møller (2003) and for $k=2$ the result can be shown by direct calculation thanks to the Palm algebra. Suppose now the statements hold for a general $k$-tuple $\underline{x}_{k}=(x_{1},\ldots,x_{k})$ of distinct points, and let $x_{k+1}\neq x_{j}$, $j=1,\ldots,k$. By the Palm algebra
$\Phi^{!}_{\underline{x}_{k+1}}=(\Phi^{!}_{\underline{x}_{k}})_{x_{k+1}}.$
Observe that $\Phi^{!}_{\underline{x}_{k}}$ has a mixture representation (by the induction hypothesis). Let $\underline{t}_{k}=(t_{1},\ldots,t_{k})\in\mathcal{T}_{k}$ be indicators describing a partition of $\underline{x}_{k}$ into $|\underline{t}|$ clusters. For ease of notation, define $\Psi_{\underline{t}_{k}}=\Phi+\sum_{h=1}^{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}$ where the $\Phi_{\zeta_{\underline{x}_{h}}}$’s are as in the statement of the theorem. Then
$\mathsf{P}_{\Phi^{!}_{\underline{x}_{k}}}=\sum_{\underline{t}_{k}\in\mathcal{T}_{k}}{\rm pr}(\underline{T}_{k}=\underline{t}_{k})\mathsf{P}_{\Psi_{\underline{t}_{k}}}.$
An application of Proposition˜S1 yields
$\mathsf{P}_{\Phi^{!}_{\underline{x}_{k+1}}}=\sum_{\underline{t}_{k}\in\mathcal{T}_{k}}w_{\underline{t}_{k}}(x_{k+1})\mathsf{P}^{!x_{k+1}}_{\Psi_{\underline{t}_{k}}},$
where $\mathsf{P}^{!x_{k+1}}_{\Psi_{\underline{t}_{k}}}$ is the law of the process $\left(\Phi+\sum_{h=1}^{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}\right)^{!}_{x_{k+1}}$ and, by Lemma˜S2,
$w_{\underline{t}_{k}}(x_{k+1})\propto{\rm pr}(\underline{T}_{k}=\underline{t}_{k})\left(\eta(x_{k+1})+\sum_{h=1}^{\underline{t}_{k}}\frac{\eta(\underline{x}_{h},x_{k+1})}{\eta(\underline{x}_{k})}\right)=:{\rm pr}(\underline{T}_{k}=\underline{t}_{k})\lambda_{\underline{t}}(x_{k+1}).$
@@
$\left(\Phi+\sum_{h=1}^{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}\right)^{!}_{x_{k+1}}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\begin{cases}\Phi+\Phi_{\zeta_{x_{k+1}}}+\sum_{h=1}^{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}&\text{with prob. }\frac{\eta(x_{k+1})}{\lambda_{\underline{t}}(x_{k+1})}\\ \Phi+\Phi_{\zeta_{(\underline{x}_{j},x_{k+1})}}+\sum_{h\neq j}{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}&\text{with prob. }\frac{\eta(\underline{x}_{h},x_{k+1})/\eta(\underline{x}_{h})}{\lambda_{\underline{t}}(x_{k+1})}\end{cases}$
@@
$\Phi^{!}_{\underline{x}_{k+1}}\stackrel{{\scriptstyle\mbox{\scriptsize d}}}{{=}}\begin{cases}\Phi+\Phi_{\zeta_{x_{k+1}}}+\sum_{h=1}^{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}&\text{with prob. }w_{\underline{t}_{k}}(x_{k+1})\frac{\eta(x_{k+1})}{\lambda_{\underline{t}}(x_{k+1})}\\ \Phi+\Phi_{\zeta_{(\underline{x}_{j},x_{k+1})}}+\sum_{h\neq j}{|\underline{t}_{k}|}\Phi_{\zeta_{\underline{x}_{h}}}&\text{with prob. }w_{\underline{t}_{k}}(x_{k+1})\frac{\eta(\underline{x}_{h},x_{k+1})/\eta(\underline{x}_{h})}{\lambda_{\underline{t}}(x_{k+1})}.\end{cases}$
The proof follows by observing that $w_{\underline{t}_{k}}(x_{k+1})\frac{\eta(x_{k+1})}{\lambda_{\underline{t}}(x_{k+1})}\propto\prod_{h=1}^{|\underline{t}_{k}|}\eta(\underline{x}_{h})\eta(x_{k+1})$ and $w_{\underline{t}_{k}}(x_{k+1})\frac{\eta(\underline{x}_{h},x_{k+1})/\eta(\underline{x}_{h})}{\lambda_{\underline{t}}(x_{k+1})}\propto\prod_{j\neq h}^{|\underline{t}|}\eta(\underline{x}_{j})\eta(\underline{x}_{h},x_{k+1})$.
S6.3 General statement and proof of Theorem˜4: the Janossy measures of shot noise Cox processes
We provide the Janossy measures for finite sncps, for any arbitrary intensity measure $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)$. Theorem˜4 corresponds to the special case with the additional assumption $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)=\rho(\mathrm{d}\gamma)G_{0}(\mathrm{d}\theta)$. This special case is discussed in the following statement as well.
@@
$j_{k}(x_{1},\ldots,x_{k})=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}(1-e^{-\gamma})\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma).$
@@
$j_{k}(x_{1},\ldots,x_{k})=k!\,{\rm pr}(\Phi(\mathbb{X})=k)\,E\left[\prod_{h=1}^{|\underline{T}|}\int_{\mathbb{X}}\prod_{j:T_{j}=h}\kappa(x_{j};\theta)G_{0}(\mathrm{d}\theta)\right],$
where the expectation is taken with respect to the indicators $\underline{T}=(T_{1},\ldots,T_{k})\in\mathcal{T}_{k}$ with distribution
${\rm pr}(\underline{T}=\underline{t})\propto\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma).$
@@
The proof follows from specializing Theorem˜S1 for a sncp. In particular, Theorem˜S1 states that
$J_{k}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k})={\rm pr}(\Phi^{!}_{(x_{1},\ldots,x_{k})}(\mathbb{X})=0)M^{(k)}_{\Phi}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}).$ (S30)
Focusing on the first term of (S30), and denoting with $\underline{x}=(x_{1},\ldots,x_{k})$,
${\rm pr}\left(\Phi^{!}_{\underline{x}}(\mathbb{X})=0\right)=E\left[{\rm pr}\left(\Phi^{!}_{\underline{x}}(\mathbb{X})=0\,|\,\underline{T}\right)\right]=E\left[{\rm pr}\left(\Phi(\mathbb{X})=0\right)\prod_{h=1}^{|\underline{T}|}{\rm pr}\left(\Phi_{\zeta_{\underline{x}_{h}}}(\mathbb{X})=0\right)\right],$ (S31)
where $\underline{T}$ and the $\Phi_{\zeta_{\underline{x}_{h}}}$’s are introduced in Theorem˜3. It follows that (S31) writes as
${\rm pr}\left(\Phi^{!}_{\underline{x}}(\mathbb{X})=0\right)$ $=E\left[{\rm pr}\left(\Phi(\mathbb{X})=0\,|\,\Lambda\right)\right]E\left[\prod_{h=1}^{|\underline{T}|}E\left[{\rm pr}\left(\Phi_{\zeta_{\underline{x}_{h}}}(\mathbb{X})=0\,|\,\zeta_{\underline{x}_{h}}\right)\right]\right]$
$=E\left[e^{-\int_{\mathbb{X}}\int_{\mathbb{X}\times\mathbb{R}_{+}}\gamma\kappa(x;\theta)\Lambda(\mathrm{d}\theta\,\mathrm{d}\gamma)\mathrm{d}x}\right]E\left[\prod_{h=1}^{|\underline{T}|}E\left[e^{-\int_{\mathbb{X}}\gamma_{\underline{x}_{h}}\kappa(x;\theta_{\underline{x}_{h}})\mathrm{d}x}\right]\right]$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}E\left[\prod_{h=1}^{|\underline{T}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\underline{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\right]$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\underline{t}\in\mathcal{T}_{k}}\left[\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\underline{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\right]{\rm pr}(\underline{T}=\underline{t}),$
where $f_{\underline{x}_{h}}$ is defined in Theorem˜3. Finally, plugging this last expression in (S30) and using Lemma˜S3 for $M^{(k)}_{\Phi}$, we obtain
$J_{k}(\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k})$ $=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\eta(\underline{x}_{h})\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}f_{\underline{x}_{h}}(\mathrm{d}\theta\,\mathrm{d}\gamma)\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k}$
$=\exp\left\{-\int_{\mathbb{X}\times\mathbb{R}_{+}}\left(1-e^{-\gamma}\right)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\right\}\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{X}\times\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)\;\mathrm{d}x_{1}\,\ldots\,\mathrm{d}x_{k},$
@@
$\begin{split}&j_{k}(x_{1},\ldots,x_{k})\\ \quad&=\exp\left\{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)\right\}\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{X}}\prod_{j:t_{j}=h}\kappa(x_{j};\theta)G_{0}(\mathrm{d}\theta)\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma).\end{split}$ (S32)
Specializing the result in Theorem˜S2 for $\nu(\mathrm{d}\theta\,\mathrm{d}\gamma)=\rho(\mathrm{d}\gamma)G_{0}(\mathrm{d}\theta)$, we obtain
${\rm pr}(N,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})=\frac{1}{N!}e^{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)}\prod_{h=1}^{C}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma),$
@@
${\rm pr}(N=k)$ $=\sum_{\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\}}{\rm pr}(N=k,\{\mathcal{C}_{1},\ldots,\mathcal{C}_{C}\})$
@@
$=\frac{1}{k!}e^{-\int_{\mathbb{R}_{+}}(1-e^{-\gamma})\rho(\mathrm{d}\gamma)}\sum_{\underline{t}\in\mathcal{T}_{k}}\prod_{h=1}^{|\underline{t}|}\int_{\mathbb{R}_{+}}e^{-\gamma}\gamma^{n_{h}}\rho(\mathrm{d}\gamma),$